OpenStudy (anonymous):

x^4+1, Find the complex zeros and write in factored form

OpenStudy (perl):

do you know the difference of squares? x^2 - y^2 = ( x - y) ( x + y )

OpenStudy (perl):

well if you dont, you know it now, makes sense so far?

OpenStudy (perl):

expand the right hand side (x-y)(x+y) = x^2 -xy + yx - y^2 = x^2 - y^2

OpenStudy (perl):

lets look at simpler case x^2 + 1 = x^2 - i^2 , do you see why? because i^2 = -1

OpenStudy (anonymous):

So (x^2-1)(x^2+1) are the complex zeroes? I thought i^2 was 1

OpenStudy (perl):

i^2 = -1, remember the definition of i, it is the square root of -1

OpenStudy (anonymous):

What would the complex zeros be?

OpenStudy (perl):

well this problem has two steps

OpenStudy (perl):

lets do a substitution. let u = x^2 then we need to factor u^2 + 1 ok so far?

OpenStudy (anonymous):

Yes

OpenStudy (perl):

and how do we factor u^2 + 1 ?

OpenStudy (perl):

ok it is true that u^2 + 1 = u^2 - i^2 STOP! do you see why, if not let me know

OpenStudy (anonymous):

I think I got that

OpenStudy (perl):

because i^2 = -1 and so u^2 +1 = u^2 - (-1) = u^2 - i^2

OpenStudy (anonymous):

Okay I get that now

OpenStudy (perl):

u^2 + 1 = u^2 - i^2 = (u - i ) ( u + i ) , by difference of squares

OpenStudy (anonymous):

Okay got that too

OpenStudy (perl):

ok now we plug back (x^2 - i ) ( x^2 + i )

OpenStudy (perl):

since u = x^2

OpenStudy (anonymous):

@perl But that would give me x^4+1 not x^4-1